# Offline Imitation of Badminton Player Behavior via Experiential Contexts and Brownian Motion

### v1

@@ III-C Experiential Context Selector (ECS)
To enable HIL in turn-based sports, we propose the ECS to leverage the experience and provide prior information to the agent, e.g., what context the rally will have in a current state. ECS leverages experience to construct a context space in which the agent is allowed to select the most relevant context as the intent of the rally, thereby ensuring that the agent’s behavior during the rally is not influenced by partially incorrect decisions from the agent.
Specifically, we first extract experiences of the current state $s_{t}^{p}$ from historical rallies by the experience extracting function $EXP$:
$\{\widetilde{\tau}_{t}^{p},{{\{\tau_{n,t}^{p}\}}\}\leftarrow EXP(s_{t}^{p},\mathcal{R})},$ (2)
@@
Figure 3: Illustration of the experience extracting function $EXP$.
The experience extraction function $EXP$ identifies historical rallies resembling a given input state, and outputs corresponding action sequences. Before training, we created a dictionary using discrete values for player, opponent, and shuttlecock positions (based on a 10x10 grid) and shot types. This dictionary efficiently stores action sequences for rallies with similar discrete states. When a new state is input, we retrieve matching rallies from the dictionary based on player, opponent, shuttlecock positions, and shot types. The intersected rallies provide experienced combinations, and their action sequences construct the context space. Figure 3 illustrates the experience extraction process. For instance, if the shuttlecock is at position 25, player at 27, opponent at 35, and shot type is clear (denoted as C), we find experiences meeting these conditions. If strict conditions are not met, we expand the range until a fitting experience is found.
Although dictionary retrieval may introduce errors due to relaxed selection criteria, the goal is to establish context spaces for the agent to learn and select appropriate contexts. Balancing the trade-off between error and generalization ability is crucial. A stricter criterion reduces errors but may limit generalization during sparse states, emphasizing the need for effective training to ensure proper context generalization.
Figure 4: Illustration of Experiential Context Selector (ECS).
Figure 4 illustrates the procedure that ECS constructs the context space and selects contexts as tactical intents based on the extracted experiences. We employ the variational autoencoders (VAE) [17, 18] architecture as the context encoder and context decoder for learning meaningful context representations from experiences. For the $t$-th step in a rally, the context encoder $q$ encodes experiences $\{\tau_{n,t}^{p}\}$ into contexts. The context encoder $q$ encodes the $n$-th experience $\tau_{n,t}^{p}$ into context $z_{n}$:
$z_{n}\sim q(\bullet|\tau_{n,t}^{p}).$ (3)
@@
$z_{ctx}=CTX(concat(\bar{z}_{ctx},x_{t}^{p})).$ (5)
As mentioned in Section III-A, there are two processes in ECS: (i) For the $t$-th step in a rally, the player process selects the player context $z_{ctx}$ in the context space constructed based on the experiences; and (ii) The target process utilizes the same context encoder $q$ to encode the correct action sequence of $\widetilde{\tau}_{t}^{p}$ to obtain the target context $\widetilde{z}_{ctx}$:
$\widetilde{z}_{ctx}\sim q(\bullet|\widetilde{\tau}_{t}^{p}).$ (6)
@@
III-D Latent Geometric Brownian Motion (LGBM)
In badminton games, players consider their opponent’s intent and next action to determine their defensive position accordingly to better receive the shuttlecock. As a result, a player’s behavior changes depending on the opponent’s intent. To enable the agent’s behavior to jointly consider the opponent’s actions, LGBM is proposed to model players as particles undergoing geometric Brownian motion in latent space, which brings the inductive bias of geometric Brownian motion into the player’s decision-making. The successful application of GBM in diverse domains, including finance systems [21] and population dynamics [22], serves as inspiration for its utilization in mitigating uncertainty within turn-based sports. As a continuous-time stochastic process with Gaussian distributed increments, GBM establishes a robust foundation, affording adaptability in mapping the latent space to diverse distributions [23]. This flexibility facilitates the capture of diverse decision distributions, accommodating a range of player strategies.
Figure 5: Illustration of Latent Geometric Brownian Motion (LGBM).
Figure 5 demonstrates the alternative simulation of geometric Brownian motion in the latent space in LGBM. This process is driven by the acquired intent from ECS and the player’s rally state. Specifically, we first concatenated the agent’s state and selected context as the agent’s position in the latent space:
$\begin{split}\widetilde{z}_{t}^{\prime}&=concat(x_{t}^{p},\widetilde{z}_{ctx}),\\ z_{t}^{\prime}&=concat(x_{t}^{p},z_{ctx}),\\ \end{split}$ (7)
@@ III-D Latent Geometric Brownian Motion (LGBM)
$\begin{split}\mathrm{\Delta}\widetilde{z}_{t}=h_{\theta}(\widetilde{z}_{t}^{\prime},t)\mathrm{\Delta}t+\sigma(\widetilde{z}_{t}^{\prime},t)\mathrm{\Delta}W_{t},\\ \mathrm{\Delta}z_{t}=h_{\phi}(z_{t}^{\prime},t)\mathrm{\Delta}t+\sigma(z_{t}^{\prime},t)\mathrm{\Delta}W_{t},\end{split}$ (10)
@@
$\begin{split}z_{t}&=z_{t-1}+\mathrm{\Delta}\widetilde{z}_{t}\enspace\mbox{(during training)},\\ z_{t}&=z_{t-1}+\mathrm{\Delta}z_{t}\enspace\mbox{(during inference)}.\end{split}$ (11)
@@
III-E Action Projection Layer
@@
$\begin{split}\langle\{\mu_{t}^{L}\},\{\sigma_{t}^{L}\},\{w_{t}^{L}\}\rangle&=W^{L}z_{t},\\ \langle\{\mu_{t}^{M}\},\{\sigma_{t}^{M}\},\{w_{t}^{M}\}\rangle&=W^{M}z_{t},\end{split}$ (13)
@@ III-F Loss Function
$\mathcal{L}=w_{1}\cdot\mathcal{L}_{pred}+w_{2}\cdot\mathcal{L}_{ctx}+w_{3}\cdot\mathcal{L}_{sde},$ (14)
@@
$\begin{split}\mathcal{L}_{pred}&=\mathcal{L}_{type}+\mathcal{L}_{land}+\mathcal{L}_{move}+\mathcal{L}_{reg},\\ \mathcal{L}_{type}&=-\sum_{T_{r}\in\{\mathcal{R}\}}\sum_{t=1}^{T}t_{t}^{p}\log\hat{t}_{t}^{p},\end{split}$ (15)
@@
$\begin{split}&\mathcal{L}_{(land,\ move)}=-\sum_{T_{r}\in\{\mathcal{R}\}}\sum_{t=1}^{T}\\ \log(&\mathcal{P}(x_{t}^{(L,\ M)},y_{t}^{(L,\ M)}|\{\mu_{t}^{(L,\ M)}\},\{\sigma_{t}^{(L,\ M)}\},\{w_{t}^{(L,\ M)}\})).\end{split}$ (16)
The regularization loss $\mathcal{L}_{reg}$ is also introduced to prevent the model from degenerating into a simple strategy by ensuring the avoidance of overlapping bivariate normal distributions, computed as the average negative distance between their means.
@@
$\mathcal{L}_{latent}=D_{KL}(\mathcal{N}(\mu_{c},\sigma_{c})||\mathcal{N}(0,1)),$ (17)
where $\mu_{c}$ and $\sigma_{c}$ are the mean and standard deviation output by the context encoder, respectively. The objective of context decoders is to reconstruct intent, which are action sequences of rallies, from selected contexts. Therefore, the reconstruction loss $\mathcal{L}_{recon}$ is the same as $\mathcal{L}_{pred}$, where we minimize the cross-entropy loss for the reconstructed shot types, the negative log-likelihood losses for the reconstructed landing and moving positions, and the average negative distance between the means of bivariate normal distributions.
The player process output needs to be close to the target process as only the player process is used during inferencing. We follow [24] to minimize the KL divergence $\mathcal{L}_{sde}$ between two SDEs in Eq. (10) of LGBM:
$\mathcal{L}_{sde}=\sum_{T_{r}\in\{\mathcal{R}\}}\sum_{t=1}^{T}\frac{1}{2}\lvert(h_{\phi}(\widetilde{z}_{t}^{\prime},t)-h_{\theta}(\widetilde{z}_{t}^{\prime},t))/\sigma(\widetilde{z}_{t}^{\prime},t)\rvert^{2}.$ (18)
IV Experiments
To assess RallyNet’s imitation performance, we conducted extensive experiments to address the alignment with the three key criteria and delved into three additional research questions, aiming to comprehensively evaluate RallyNet’s effectiveness and its value to the badminton community:
@@ IV Experiments
    Rally Duration Realism: Does RallyNet accurately replicate the length of rallies, providing a realistic representation of the duration observed in real matches?
@@
    Outcome Consistency: Are the outcomes of rallies generated by RallyNet consistent with the real-world results? Specifically, does the win rate of rallies simulated by the learned agent align with the win rate of real players?
@@
    Ablation Studies: Does each component of RallyNet affect the prediction results?
  • RQ5
    Sensitivity Analysis: How sensitive is RallyNet’s performance to variations in input parameters and conditions?
  • RQ6
@@
IV-A Experimental Setup
IV-A1 Badminton Dataset
@@
RallyNet is evaluated on the largest badminton singles dataset, which is the only publicly available benchmark in turn-based sports [9, 25]. This dataset comprises 75 singles matches played by 31 players from 2018 to 2021, totaling 180 sets, 4,325 rallies, and 43,191 strokes. To construct the action space, we used the 12 shot types defined by [26], namely receiving, short service, long service, net shot, clear, push/rush, smash, defensive shot, drive, lob, drop, and can’t reach. The range of each dimension of the badminton court was rescaled to [-1,1] to ensure that the model would not be affected by different ranges. We trained our model on the first 80% of the rallies, and the remaining 20% of rallies were used to evaluate the performance of our model as well as the 5-fold cross-validation for tuning the hyper-parameters.
IV-A2 Baselines
We compared RallyNet against several baselines, including:
  • Random agent, which samples the shot type as well as both the landing and moving positions uniformly randomly.
  • Rule-based agent, which samples from the distribution of the shot type and from both landing and moving distributions which are obtained through the experience extracting function of RallyNet.
  • IQ-Learn [27], a state-of-the-art model-free offline inverse reinforcement learning algorithm.
  • Behavior Cloning (BC) [28], which learns a mapping from a state to an action.
  • Hierarchical Behavioral Cloning (HBC) [29], which learns an options-type hierarchical policy from expert demonstrations.
  • ShuttleNet [9], which is a specialized model designed for stroke forecasting. It fuses the contexts of rally progress and player styles to predict the shot type and the landing position based on past information. It is noted that ShuttleNet requires at least two steps to encode players’ contexts; therefore, it cannot be tested only on the initial state. Since ShuttleNet originally predicted only the shot type and landing positions, we extended its capabilities by incorporating a module similar to the landing position prediction to also enable the prediction of moving positions.
  • DyMF [25], which is a specialized model designed for badminton movement forecasting. It employs interaction style extractors to adeptly capture nuanced player interactions and evolving tactics across rallies. DyMF predicts players’ shot types and movements. Given that a player’s receiving position essentially determines the landing spot, DyMF seamlessly aligns with our problem setting. However, akin to ShuttleNet, DyMF relies on at least two steps of past information for future predictions, limiting its capacity to recover the entire rally from an initial state.
IV-A3 Implementation Details
To avoid the agent’s predictions for landing positions and movement positions encompassing the entire court and degenerating the strategy, we imposed a constraint on the standard deviation of the bivariate Gaussian distribution for our method and all baselines. The standard deviation was limited to the range of [-0.1, 0.1]. To improve the model’s generalization capability and to mitigate over-reliance on experiences for decision-making, we utilized an experience extracting function that restricted the number of experiences for the current state to 5. This was done because a substantial amount of experience cannot be guaranteed during testing, particularly in the later stages of a turn-based setting when the states become sparse.
The dimension of state embedding was set to 80, and the dimension of context embedding was set to 128. The batch size was 32 and the number of training epochs was 100 using Adam optimizer. The learning rate was set to 0.0001. For HBC implementations, the number of options for HBC was treated as a hyperparameter, with experiments conducted using various numbers of options 2, 4, 8, 16. The optimal result was observed when the number of options was set to 4. All training and evaluation experiments were performed on a machine equipped with an Intel i7-8700 3.2GHz CPU, Nvidia GTX 3060 12GB GPU, and 32GB RAM.
@@ IV-A3 Implementation Details
TABLE I: Quantitative results. The best is in boldface and the second best is underlined. Since ShuttleNet and DyMF predict based on past information, the symbol -- denotes that the result is unavailable.
@@
IV-A4 Evaluation Metrics
Given the absence of prior research on the imitation of behaviors in turn-based sports, we proposed 4 evaluation metrics to measure the similarity between generated rallies and true player rallies. To evaluate the results of shot type prediction, we used Connectionist Temporal Classification (CTC) loss [30] for uncertainty measurement, which is defined as the negative log-likelihood of the labels given input sequences. To evaluate the results of the predicted landing and moving positions, we used Dynamic Time Warping (DTW) [31] to calculate the distance between generated position sequence and the true position sequence, which can assess the similarity of two sequences at a global level, as used recently in [32, 33]. Using the DTW distance and the CTC loss has the benefit that the similarity of two sequences can be assessed even if each sequence is of a different length. Moreover, inspired by [34, 35], we further provide an overall comparison of the algorithms by introducing a metric termed Rule-based agent Normalized Score (RNS) defined as:
$RNS=\frac{(Random_{score}-Agent_{score})}{(Random_{score}-Rule_{score})},$ (19)
where the subscript $score$ indicates the metric used and can be either DTW distance or CTC loss, and $Agent_{score}$ is the performance of the agent. $Random_{score}$ and $Rule_{score}$ are the performance of the random agent and the rule-based agent, respectively.
To provide an overall comparison of the algorithms, we present the Mean RNS (MRNS), which is defined as the average over the RNS values under the three base metrics. This metrics selection aligns with our emphasis on tactical execution, a domain where conventional metrics like Mean Squared Error (MSE) or Cross-Entropy (CE) used in [9, 25] fail to reflect the intricate similarities inherent in decision sequences. This limitation becomes evident when minor sequence deviations arise. For instance, even if a strategy compels opponents to traverse the court substantially, a minor delay by a single stroke should still be recognized as similar. All the results are the average of 5 different random seeds.
IV-B Quantitative Results (RQ1)

### v2

@@ 4.1 Experiential Context Selector (ECS)
To enable HIL in turn-based sports, we propose the ECS to leverage the experience and provide prior information to the agent, e.g., what context the rally will have in a current state. ECS leverages experience to construct a context space in which the agent is allowed to select the most relevant context as the intent of the rally, thereby ensuring that the agent’s behavior during the rally is not influenced by partially incorrect decisions from the agent. Specifically, we first extract experiences of the current state $s_{t}^{p}$ from historical rallies by the experience extracting function $EXP$:
$\small\{\widetilde{\tau}_{t}^{p},{{\{\tau_{n,t}^{p}\}}\}\leftarrow EXP(s_{t}^{p},\mathcal{R})},$ (1)
@@
We employ the variational autoencoders (VAE) [15, 24] architecture as the context encoder and context decoder for learning meaningful context representations from experiences. For the $t$-th step in a rally, the context encoder $q$ encodes experiences $\{\tau_{n,t}^{p}\}$ into contexts. The context encoder $q$ encodes the $n$-th experience $\tau_{n,t}^{p}$ into context $z_{n}$:
$\small z_{n}\sim q(\bullet|\tau_{n,t}^{p}).$ (2)
@@
$\small z_{ctx}=CTX(concat(\bar{z}_{ctx},x_{t}^{p})).$ (4)
There are two processes in ECS: (i) For the $t$-th step in a rally, the player process selects the player context $z_{ctx}$ in the context space constructed based on the experiences; and (ii) The target process utilizes the same context encoder $q$ to encode the correct action sequence of $\widetilde{\tau}_{t}^{p}$ to obtain the target context $\widetilde{z}_{ctx}$:
$\small\widetilde{z}_{ctx}\sim q(\bullet|\widetilde{\tau}_{t}^{p}).$ (5)
@@
4.2 Latent Geometric Brownian Motion (LGBM)
In badminton games, players consider their opponent’s intent and next action to determine their defensive position accordingly to better receive the shuttlecock. As a result, a player’s behavior changes depending on the opponent’s intent. To enable the agent’s behavior to jointly consider the opponent’s actions, LGBM is proposed to model players as particles undergoing geometric Brownian motion in latent space, which brings the inductive bias of geometric Brownian motion into the player’s decision-making. The flexibility of GBM in capturing diverse decision distributions has led to its successful application in various domains, such as finance systems [22] and population dynamics [7]. These successes serve as inspiration for capturing the inherent uncertainty and interaction in badminton.
Specifically, we first concatenated the agent’s state and selected context as the agent’s position in the latent space:
$\small\begin{split}\widetilde{z}_{t}^{\prime}=concat(x_{t}^{p},\widetilde{z}_{ctx}),\quad z_{t}^{\prime}=concat(x_{t}^{p},z_{ctx}),\\ \end{split}$ (6)
@@ 4.2 Latent Geometric Brownian Motion (LGBM)
$\small\begin{split}\mathrm{\Delta}\widetilde{z}_{t}=h_{\theta}(\widetilde{z}_{t}^{\prime},t)\mathrm{\Delta}t+\sigma(\widetilde{z}_{t}^{\prime},t)\mathrm{\Delta}W_{t},\quad\mathrm{\Delta}z_{t}=h_{\phi}(z_{t}^{\prime},t)\mathrm{\Delta}t+\sigma(z_{t}^{\prime},t)\mathrm{\Delta}W_{t},\end{split}$ (9)
@@
$\small\begin{split}z_{t}=z_{t-1}+\mathrm{\Delta}\widetilde{z}_{t}\enspace\mbox{(during training)},\quad z_{t}=z_{t-1}+\mathrm{\Delta}z_{t}\enspace\mbox{(during inference)}.\end{split}$ (10)
@@
4.3 Action Projection Layer
@@
$\small\begin{split}\langle\{\mu_{t}^{L}\},\{\sigma_{t}^{L}\},\{w_{t}^{L}\}\rangle=W^{L}z_{t},\quad\langle\{\mu_{t}^{M}\},\{\sigma_{t}^{M}\},\{w_{t}^{M}\}\rangle=W^{M}z_{t},\end{split}$ (12)
@@ 4.4 Loss Function
$\small\mathcal{L}=w_{1}\cdot\mathcal{L}_{pred}+w_{2}\cdot\mathcal{L}_{ctx}+w_{3}\cdot\mathcal{L}_{sde},$ (13)
@@
$\small\begin{split}\mathcal{L}_{pred}=\mathcal{L}_{type}+\mathcal{L}_{land}+\mathcal{L}_{move}+\mathcal{L}_{reg},\quad\mathcal{L}_{type}=-\sum_{T_{r}\in\{\mathcal{R}\}}\sum_{t=1}^{T}t_{t}^{p}\log\hat{t}_{t}^{p},\end{split}$ (14)
@@
$\small\begin{split}\mathcal{L}_{(land,\ move)}=-\sum_{T_{r}\in\{\mathcal{R}\}}\sum_{t=1}^{T}\log(\mathcal{P}(x_{t}^{(L,\ M)},y_{t}^{(L,\ M)}|\{\mu_{t}^{(L,\ M)}\},\{\sigma_{t}^{(L,\ M)}\},\{w_{t}^{(L,\ M)}\})).\end{split}$ (15)
The regularization loss $\mathcal{L}_{reg}$ is introduced to prevent the model from degenerating into a simple strategy by ensuring the avoidance of overlapping distributions, computed as the average negative distance between their means.
@@
$\small\mathcal{L}_{latent}=D_{KL}(\mathcal{N}(\mu_{c},\sigma_{c})||\mathcal{N}(0,1)),$ (16)
where $\mu_{c}$ and $\sigma_{c}$ are the mean and standard deviation output by the context encoder, respectively. The objective of context decoders is to reconstruct intent, which are action sequences of rallies, from selected contexts. Therefore, the reconstruction loss $\mathcal{L}_{recon}$ is the same as $\mathcal{L}_{pred}$. Finally, we follow [18] to minimize the KL divergence $\mathcal{L}_{sde}$ between two SDEs in Eq. (9) of LGBM:
$\small\mathcal{L}_{sde}=\sum_{T_{r}\in\{\mathcal{R}\}}\sum_{t=1}^{T}\frac{1}{2}\lvert(h_{\phi}(\widetilde{z}_{t}^{\prime},t)-h_{\theta}(\widetilde{z}_{t}^{\prime},t))/\sigma(\widetilde{z}_{t}^{\prime},t)\rvert^{2}.$ (17)
5 Experiments
To evaluate RallyNet’s imitation performance, we conducted comprehensive experiments addressing three key criteria and explored case studies to gauge its value to the badminton community [^1] Our code and dataset are already publicly available as the GitHub repository https://github.com/wywyWang/CoachAI-Projects.:
@@ 5 Experiments
    Rally Duration Realism: Does RallyNet replicate the length of rallies, providing a realistic representation of the duration observed in real matches?
@@
    Outcome Consistency: Are the outcomes generated by RallyNet consistent with the real-world results? Specifically, does the win rate of rallies simulated by the learned agent align with the win rate of real players?
@@
5.1 Experimental Setup
@@
Badminton Dataset. RallyNet is evaluated on the largest badminton singles dataset, which is the only publicly available benchmark in turn-based sports [29, 4]. This dataset comprises 75 singles matches played by 31 players from 2018 to 2021, totaling 180 sets, 4,325 rallies, and 43,191 strokes. To construct the action space, we used the 12 shot types defined by [27], namely receiving, short service, long service, net shot, clear, push/rush, smash, defensive shot, drive, lob, drop, and can’t reach. For each match, the 80% of rallies are designated as training data, while the remaining 20% are reserved for testing, ensuring comprehensive integration of each player’s historical data into the model. We utilize 5-fold cross-validation to fine-tune hyperparameters.
Baselines. We compare RallyNet against several baselines, including: 1) Random agent, which samples the shot type as well as both the landing and moving positions uniformly randomly; 2) Rule-based agent, which samples from the distribution of the shot type and from both landing and moving distributions which are obtained through the experience extracting function; 3) IQ-Learn [8], a state-of-the-art model-free offline IRL; 4) Behavior Cloning (BC) [20]; 5) Hierarchical Behavioral Cloning (HBC) [31], which learns an options-type hierarchical policy from demonstrations; 6) ShuttleNet [29], which is a specialized sequence-to-sequence model designed for stroke forecasting. It fuses the contexts of rally progress and player styles to predict the shot type and the landing position based on past information. It is noted that ShuttleNet is not designed for imitation; requiring at least two steps to encode players’ contexts, it cannot be tested solely from the initial state. 7) DyMF [4], which is a specialized sequence-to-sequence model tailored for badminton movement forecasting. It employs dynamic graphs and hierarchical fusion to adeptly capture player interactions. Similar to ShuttleNet, DyMF relies on past information for future predictions, limiting its ability to recover the entire rally from an initial state.
Evaluation Metrics. Given the absence of prior research on the imitation of behaviors in turn-based sports, we proposed 4 evaluation metrics to measure the similarity between generated rallies and true player rallies. To evaluate the results of shot type prediction, we used Connectionist Temporal Classification (CTC) loss [12] for uncertainty measurement, which is defined as the negative log-likelihood of the labels given input sequences. To evaluate the results of the predicted landing and moving positions, we used Dynamic Time Warping (DTW) [2] to calculate the distance between generated position sequence and the true position sequence, which can assess the similarity of two sequences at a global level, as used recently in [26]. Using the DTW distance and the CTC loss has the benefit that the similarity of two sequences can be assessed even if each sequence is of a different length. Moreover, inspired by [1], we further provide an overall comparison of the algorithms by introducing a metric termed Rule-based agent Normalized Score (RNS) defined as: $RNS=\frac{(Random_{score}-Agent_{score})}{(Random_{score}-Rule_{score})},$ where the subscript $score$ indicates the metric used and can be either DTW distance or CTC loss, and $Agent_{score}$ is the performance of the agent. $Random_{score}$ and $Rule_{score}$ are the performance of the random agent and the rule-based agent, respectively. To provide an overall comparison, we present the Mean RNS (MRNS), which is defined as the average over the RNS scores under the 3 metrics. This metrics aligns with our emphasis on tactical execution, where Mean Squared Error (MSE) or Cross-Entropy (CE) used in [29, 4] might overlook similarities within decision sequences. For instance, if a strategy compels opponents to traverse the court substantially, a minor delay by a single stroke should still be recognized as similar. All the results are the average of 5 different random seeds.
@@
Table 1: Quantitative results. The best is in boldface and the second best is underlined. Since ShuttleNet and DyMF predict based on past information, the symbol -- denotes that the result is unavailable.
@@
5.2 Quantitative Results (RQ1)